OpenStudy (anonymous):

Case worked for $6/hr one week and $7/hr the next week testing video games. He worked 5 more hours the second week that the first week. He earned a total of $347. Hom many hours did he work each week?

OpenStudy (anonymous):

Okay, so we can use "x" as the number of hours. x*6 is what he earned the first week (x+5)*7 is what he earned the second week, since he did 5 more hours. x*6 + (x+5)*7 = 347 6x + 7x + 35 = 347 13x = 347 - 35 13x = 312 x = 24 So now we know x is 24. And we know the second week, it was x+5, so the first week he worked 24 hours and the second week, 29.